Suppose that an industry is dominated BT a single producer, the demand for it's product is Q=300-6p.suppose further that TC=5Q+1/480Q×2 is the cost function of the firm


1
Expert's answer
2021-04-13T09:36:38-0400

The profit-maximizing quantity is produced at MR = MC, so:

Q = 300 - 6p or p = 50 - 1/6Q,

"MR = TR'(Q) = 50 - 1\/3Q,"

"MC = TC'(Q) = 5 + 1\/240Q,"

50 - 1/3Q = 5 + 1/240Q,

81/240Q = 45,

Q = 45×240/81 = 133.33 units.

p = 50 - 1/6×133.33 = 27.78.